OpenStudy (anonymous):

Tell whether the lines for each pair of equations are parallel, perpendicular, or neither. y=-4x+3 -2x=8y=5 A. Parallel B. Perpendicular C. Neither Tell whether the lines for each pair of equations are parallel, perpendicular, or neither. y=-1/4x+10 -2x+8y=6 A. parallel B. Perpendicular C. Neither Write y=(-5/8)x+3 in standard form using integers. A. 5x+8y=-24 B. 5x+8y=24 C. 5x-8y=24 D.-5x+8y=24 PLEASE HELP!!!

OpenStudy (anonymous):

@ganeshie8 @agent0smith @Ashleyisakitty

OpenStudy (anonymous):

@robtobey @Nurali @phi

OpenStudy (anonymous):

express all lines in y = mx+ c format m gives the slope of the line for 2 parallel lines slopes are equal ie. m1=m2 for perpendicular lines m1.m2 = -1

OpenStudy (anonymous):

But that doesn't help me with the questions.

OpenStudy (anonymous):

why ? showed you how to do such problems ... try yourself now.

OpenStudy (anonymous):

I don't know how to do it!

OpenStudy (anonymous):

first one is neither , second one is parallel , third answer : 5x+8y=24

OpenStudy (anonymous):

can you help me with one last one?

OpenStudy (anonymous):

y=(-5/8)x+3 so you need to remove the 8 from the denominator you can do this by multiplying 8 on both sides of the equation 8y = 8(-5/8) x + 8x3 or 8y = -5x + 24 or 8y+5y = 24

OpenStudy (anonymous):

No, I mean another one?

OpenStudy (anonymous):

what is the prob ?

OpenStudy (anonymous):

write the equation of a line that is perpendicular to the given line and that passes through the given points. y-2=7/3(x=5); (-4,9) A. y-9= -3/7(x-4) B. y-9=3/7(x-4) C. y-9=-3/7(x+4) D. y-9=3/7(x+4)

OpenStudy (anonymous):

y-2=7/3(x=5); check this part .. (x=5) should either be x+5 or x-5

OpenStudy (anonymous):

I don't understand

OpenStudy (anonymous):

an equality has two sides with one equal to sign in between but your problem has y-2=7/3(x=5) two equal to signs ... which means the sum is incorrect

OpenStudy (anonymous):

so it is none of the answers?

OpenStudy (anonymous):

yes